OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!! Justin starts running south at 8:00 AM at a rate of 6 minutes per mile. Justin then starts running west at a rate of 10 minutes per mile. Justin stops running at 8:50 AM. If Justin runs a total of 6 miles, how many miles did he run in a westerly direction?

OpenStudy (anonymous):

justin bieber?

OpenStudy (anonymous):

No, seriously, help please)

OpenStudy (dumbcow):

break it into 2 parts, d1 and d2 , t1 and t2 t1 = 6d1 t2 = 10d2 t1 + t2 = 50 d1 + d2 = 6 substitute 6d1 +10d2 = 50 d1 + d2 = 6 solve system of equations

OpenStudy (anonymous):

thank you))

OpenStudy (dumbcow):

welcome

OpenStudy (anonymous):

no thank you for me :(